Let A and B be n × n matrices. Prove that trAB = trBA and trA = trAt

.


1
Expert's answer
2021-07-22T09:14:30-0400

A and B are matrices of order "n" .

We know trace is sum of diagonal elements of the matrix.

We need to prove "tr" ("AB" ) = "tr" ("BA" )


When we multiply matrix A with B i.e AB or when we multiply matrix B with A i.e BA in both cases elements in the diagonal will remain the same. So sum of diagonal elements in AB =sum of diagonal elements in BA .

So "tr" ("AB" ) = "tr" ("BA" ) (proved)


Now we need to prove "tr" ("A^t)=tr(A)"

So in "A^t" diagonal elements will remain same as that of A .So sum of diagonal elements in A=sum of diagonal elements in "A^t"

"\implies tr(A^t)=tr(A)" (proved)